OpenStudy (anonymous):

Find the solution for 3-2x=2-x

OpenStudy (anonymous):

3-2x=2-x 3-2=-x+2x 1=x is the answer.

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

3-2x = 2-x take x across to LHS -x = 2 -3 x = 1